# Formulation of entropy-conservative discretizations for compressible flows of thermally perfect gases

### v2

@@ 4 Numerical results
Numerical tests are presented to assess the EC properties and the robustness of the proposed formulation; all simulations have been carried out using the dimensionless form of the equations. Time integration has been performed with a standard fourth-order explicit Runge–Kutta method, with a sufficiently small CFL number to limit time-integration errors. For the thermodynamic framework, we have adopted the experimental fit for $c_{v}(T)$ given by the five-coefficient approximation reported in [42], where the only nonzero coefficients are $c_{-2},\,c_{0},\,c_{1},\,c_{2},\,c_{3}$. In the following results, the specific gas chosen is high-enthalpy methane, whose heat capacities strongly depend on temperature [42, 43]. Moreover, the nonpolarity of such molecule allows van der Waals forces to be neglected, especially at high temperatures where its modelling as a thermally perfect gas is preferable. Regarding the spatial discretization, we use the novel formulation (11)–(12) (together with the numerical flux for momentum given in Eq. (2)), here termed AEC-TP(N), with a sufficiently high value of $N$; in all our tests we checked that the value of $N$ guaranteeing the convergence to the EC scheme in Eq. (10) is typically not greater than $5$. The novel discretization is compared to the one proposed by Gouasmi et al., reported in Eq. (14) and generalized to include also negative exponents in the expression of the specific heat capacities. Additionally, results are provided for other widely used discretizations, such as the formulation proposed by Ranocha [23]
@@
The Ranocha’s flux in Eq. (21) is exactly EC and PEP for calorically perfect gases, and is also KEP. The KEEP and JP fluxes in Eq. (22) and (23) are KEP, but not EC, although the KEEP formulation has noticeable entropy conservation properties for calorically perfect gases [17, 27, 26]. In order to better evaluate the effect that the discretizations have on the evolution of quantities such as entropy and kinetic energy, we define a normalized difference of the integral value with respect to its value at time zero. For a generic variable $\phi$, that is
@@ 4.1 Inviscid doubly periodic jet flow
In this section we report the case of the inviscid doubly periodic jet flow. The flow is solved in the periodic domain $(x,y)\in[0,L]\times[-L/4,L/4]$, with $L=1$, on a coarse mesh consisting of $N_{P}=65\times 33$ evenly spaced points by means of second-order, central, spatial discretizations. CFL is set equal to $0.01$. The initial conditions are chosen as it follows
$\begin{dcases}u(x,y^{\pm},0)&=A_{u}\left[1\mp B_{u}\tanh{\dfrac{y\pm L/10}{\delta}}\right]\\ v(x,y,0)&=\epsilon\sin{(2n\pi x/L)}\\ T(x,y^{\pm},0)&=A_{T}\left[\dfrac{3}{2}\pm B_{T}\tanh{\dfrac{y\pm L/10}{\delta}}\right]\\ p(x,y,0)&=p_{0}\end{dcases}$ (24)
@@
The double-jet flow has been chosen because it displays a wide temperature range, which makes the hypothesis of calorically perfect gas too stringent—especially for chemical species such as methane. Second-order spatial discretization is used here to test the basic versions of the fluxes as discussed in Sec. 2.1; high-order formulations are used in the next test case in Sec. 4.2. Fig. 1(a) shows the asymptotic convergence of the AEC-TP(N) formulation with the increasing number of expansion terms. For this particular test case, a number $N=5$ of terms in the asymptotic expansion is sufficient to reach machine-zero entropy production, indicating convergence to the exact EC scheme in Eq. (10). In Fig. 1(b) the error on entropy conservation is shown for the various numerical schemes here considered. Both the AEC-TP(5) and the fluxes proposed by Gouasmi et al. show exact entropy conservation, while the discretizations designed for ideal gases reveal an error which accumulates in time. The same applies to the AEC-TP(0) which, however, shows a better behaviour for this particular test case.
@@ 4.2 Inviscid three-dimensional Taylor–Green Vortex
Figure 2: Inviscid three-dimensional Taylor–Green vortex at $\operatorname{\mathit{M}}=0.1$. (a) entropy-conservation error (b) kinetic energy evolution (c) internal energy evolution (d) kinetic energy-related pressure term evolution. For graphical clarity, data are sampled every 30 time steps for figures (a)-(c) and every 500 time steps for figure (d).
@@
Fig. 2(a) shows the entropy-conservation properties of the desired schemes. As already observed in [39], the TGV produces “well-behaved" entropy evolutions regardless of the gas model if the numerical scheme is, even approximately, EC for ideal gases. We could attribute such behaviour to the test case itself, which provides relatively small variations of the thermodynamic quantities. However, exact conservation of entropy for the AEC-TP(N) with $N=3$ (which is the value guaranteeing convergence for this test case) and for the Gouasmi scheme is evident in Fig. 2(a), where the nondimensionalized error on entropy conservation is zero to machine precision. The KEEP and Ranocha formulations display entropy production which is approximately three orders of magnitude greater, whereas the JP scheme completely fails to keep the entropy production under control. Despite their almost identical behaviour on entropy conservation, the AEC-TP and Gouasmi formulations show important discrepancies when evaluated on other observables. An example is given by the evolution of global kinetic energy, which is reported in Fig. 2(b). Although total kinetic energy is not conserved in general in compressible inviscid flows, for the inviscid TGV test under consideration it is expected that global kinetic energy does not vary in time, because of the periodicity [18]. The Gouasmi formulation is designed to be KEP, as all the other formulations we are considering, in the sense that the discretization of the advective terms does not influence the kinetic energy balance. However, as documented for the Chandrashekar formulation in [18], the discretization of the pressure term of the momentum equation as $R\overline{\rho}\,/(\overline{1/T}\,)$ produces nonnegligible kinetic-energy dissipation with respect to all of the other schemes, which enforce the discretization of the pressure flux as $\overline{p}\,$. Fig. 2(c) shows the evolution of global internal energy, with an analogous production for the Gouasmi scheme due to the transfer of energy allowed by the pressure-work terms. The effect of the spatially integrated pressure work contribution is shown in Fig. 2(d), where the integrated time derivative of kinetic energy is shown as representative of the nonconservative pressure term in the discrete kinetic energy evolution equation, since the convective term is in locally conservative form. The nonzero average displayed by the Gouasmi scheme reflects its failure in keeping the global kinetic energy constant.
@@
An additional difference between the novel and previous formulations is seen in the fluctuations of the thermodynamic quantities of the density $\rho^{\prime}_{\mathrm{rms}}$ and temperature $T^{\prime}_{\mathrm{rms}}$, reported in Fig. 3, in which $\phi^{\prime}_{\mathrm{rms}}=\sqrt{\frac{1}{N_{p}}\sum_{i=1}^{N}(\phi_{i}-\phi_{\mu})^{2}}$ and $\phi_{\mu}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\phi_{i}$ is the average value of $\phi$ in the domain. In fact, as it has been already shown for the JP scheme ([12, 17]), it seems that energy fluxes whose pressure term is not built as in (2) produce fluctuations which do not reach a steady-state after the thermalization of the flow field, which is the expected result for homogeneous isotropic turbulence calculations [47].
@@ 5 Conclusions
Starting from the EC condition valid for a generic real gas, we demonstrated how the specific structure of the EoS for perfect gases can be exploited to eliminate singularities in the formulation of the numerical flux, thereby simplifying the resulting expressions. The only remaining singularity arises from the presence of logarithmic means, which can be addressed using the algorithm by Ismail and Roe [22] or by constructing a hierarchy of asymptotically entropy-conserving schemes thorough a Taylor series expansion, as in the case of calorically perfect gases. Importantly, this result is independent of the specific thermally perfect model used to relate internal energy and absolute temperature, supporting its use in a wide range of applications. As illustrative examples, we have considered both a polynomial-based model for the isochoric specific heat and the Rigid-Rotor Harmonic-Oscillator model, each leading to a different variant of the EC scheme.
@@
The results presented in this paper open several avenues for future development. One natural extension is the construction of entropy-stable schemes by adding appropriate dissipation terms to the newly introduced EC fluxes, providing a desirable property for the development of robust high-order schemes [48]. This extension would allow a physically accurate evolution of kinetic energy together with a thermodynamically consistent inclusion of dissipation, making the schemes suitable for under-resolved or shock-dominated flows where entropy conservation is no longer physically justified.
Moreover, while the current formulation is restricted to nonreacting gas mixtures, the generality of the approach suggests potential extensions to reacting flows, including chemically reactive mixtures and multitemperature models.
@@ Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. We thank Dr. Ayaboe K. Edoh for reading an early draft of the manuscript and providing thoughtful feedback.

### v3

@@ 4 Numerical results
Numerical tests are presented to assess the EC properties and the robustness of the proposed formulation; all simulations have been carried out using the dimensionless form of the equations. Time integration has been performed with a standard fourth-order explicit Runge–Kutta method, with a sufficiently small CFL number to limit time-integration errors. For the thermodynamic framework, we have adopted the experimental fit for $c_{v}(T)$ given by the five-coefficient approximation reported in [48], where the only nonzero coefficients are $c_{-2},\,c_{0},\,c_{1},\,c_{2},\,c_{3}$, whose values are reported in Tab. 1 for convenience.
Table 1: NASA 5-coefficient polynomial for CH4 (Low temperature range), [48].
Temperature range (K) 2981300298-1300
$c_{-2}$ 0.678565
$c_{0}$ -0.703029
$c_{1}$ 108.4773
$c_{2}$ -42.52157
$c_{3}$ 5.862788
In the dimensionless framework $R=1$, thus the equation of state reads $p=\rho T$ and the Mayer’s relation is $c_{p}-c_{v}=1$. Reference energy, which is used to nondimensionalize internal energy as well as enthalpy, is $E_{\mathrm{r}}^{*}=R^{*}T_{\mathrm{r}}^{*}$, with the asterisk indicating dimensional quantities. In the following results, the specific gas chosen is high-enthalpy methane, whose heat capacities strongly depend on temperature [48, 49]. Moreover, the nonpolarity of such molecule allows van der Waals forces to be neglected, especially at high temperatures where its modelling as a thermally perfect gas is preferable. Regarding the spatial discretization, we use the novel formulation (11)–(12) (together with the numerical flux for momentum given in Eq. (2)), here termed AEC-TP(N), with a sufficiently high value of $N$; in all our tests we checked that the value of $N$ guaranteeing the convergence to the EC scheme in Eq. (10) is typically not greater than $5$. The novel discretization is compared to the one proposed by Gouasmi et al., reported in Eq. (14) and generalized to include also negative exponents in the expression of the specific heat capacities. Additionally, results are provided for other widely used discretizations, such as the formulation proposed by Ranocha [24]
@@
The Ranocha’s flux in Eq. (21) is exactly EC for calorically perfect gases, and is also KEP. The KEEP and JP fluxes in Eq. (22) and (23) are KEP, but not EC, although the KEEP formulation has noticeable entropy conservation properties for calorically perfect gases [18, 34, 33]. In order to better evaluate the effect that the discretizations have on the evolution of quantities such as entropy and kinetic energy, we define a normalized difference of the integral value with respect to its value at time zero. For a generic variable $\phi$, that is
@@ 4.1 Inviscid doubly periodic jet flow
In this section we report the case of the inviscid doubly periodic jet flow. The flow is solved in the periodic domain $(x,y)\in[0,L]\times[-L/4,L/4]$, with $L=1$, on a coarse mesh consisting of $N_{P}=65\times 33$ evenly spaced points by means of second-order, central, spatial discretizations. CFL is set equal to $0.01$. The initial conditions are chosen as follows:
$\begin{dcases}u(x,y^{\pm},0)&=A_{u}\left[1\mp B_{u}\tanh{\dfrac{y\mp L/10}{\delta}}\right]\\ v(x,y,0)&=\epsilon\sin{(2n\pi x/L)}\\ T(x,y^{\pm},0)&=A_{T}\left[\dfrac{3}{2}\pm B_{T}\tanh{\dfrac{y\mp L/10}{\delta}}\right]\\ p(x,y,0)&=p_{0}\end{dcases}$ (24)
@@
The double-jet flow has been chosen because it displays a wide temperature range, which makes the hypothesis of calorically perfect gas too stringent—especially for chemical species such as methane. Second-order spatial discretization is used here to test the basic versions of the fluxes as discussed in Section 2.1; high-order formulations are used in the next test case in Section 4.2.
Fig. 1 reports instantaneous visualizations of the entropy and pressure fields for illustrative purposes, carried out with the AEC-TP(5) discretization at a higher spatial resolution ($257\times 129$). Entropy is shown in Fig. 1(a) and it remarks the roll-up behaviour provided by the transversal jet. As expected, pressure variations (shown in Fig. 1(b), appearing as compression/expansion zones arising from the rolled-up regions) are relatively small if compared to thermal ones, which represent the major contribution to the entropy field.
Figure 1: Instantaneous visualization of the inviscid doubly periodic jet flow at $t/t_{c}=4$. Results have been obtained on a refined grid consisting in $N_{x}\times N_{y}=257\times 129$ evenly spaced points to guarantee a satisfactory visualization of the main flow-field structures. Here, CFL has been set to $0.5$.
Fig. 2(a) shows the asymptotic convergence of the AEC-TP(N) formulation with the increasing number of expansion terms. For this particular test case, a number $N=5$ of terms in the asymptotic expansion is sufficient to reach machine-zero entropy production, indicating convergence to the exact EC-TP scheme in Eq. (10). In Fig. 2(b) the error on entropy conservation is shown for the various numerical schemes here considered. Both the AEC-TP(5) and the fluxes proposed by Gouasmi et al. show exact entropy conservation, while the discretizations designed for ideal gases reveal an error which accumulates in time. The same applies to the AEC-TP(0) which, however, shows a better behaviour for this particular test case.
@@ 4.2 Inviscid three-dimensional Taylor–Green Vortex
Figure 3: Inviscid three-dimensional Taylor–Green vortex at $\operatorname{\mathit{M}}=0.1$. (a) entropy-conservation error (b) kinetic energy evolution (c) internal energy evolution (d) kinetic energy-related pressure term evolution. For graphical clarity, data are sampled every 30 time steps for Figs. (a)-(c) and every 500 time steps for Fig. (d).
@@
Fig. 3(a) shows the entropy-conservation properties of the considered schemes. As already observed in [45], the TGV produces “well-behaved" entropy evolutions regardless of the gas model if the numerical scheme is, even approximately, EC for ideal gases. We could attribute such behaviour to the test case itself, which provides relatively small variations of the thermodynamic quantities. However, exact conservation of entropy for the AEC-TP(N) with $N=3$ (which is the value guaranteeing convergence for this test case) and for the Gouasmi scheme is evident in Fig. 3(a), where the nondimensionalized error on entropy conservation is zero to machine precision. The KEEP and Ranocha formulations display entropy production which is approximately three orders of magnitude greater, whereas the JP scheme completely fails to keep the entropy production under control. Despite their almost identical behaviour on entropy conservation, the AEC-TP and Gouasmi formulations show important discrepancies when evaluated on other observables. An example is given by the evolution of global kinetic energy, which is reported in Fig. 3(b). Although total kinetic energy is not conserved in general in compressible inviscid flows, for the inviscid TGV test under consideration it is expected that global kinetic energy does not significantly vary in time [19]. The Gouasmi formulation is designed to be KEP, as all the other formulations we are considering, in the sense that the discretization of the advective terms does not influence the kinetic energy balance. However, as documented for the Chandrashekar formulation in [19], the discretization of the pressure term of the momentum equation as $R\overline{\rho}\,/(\overline{1/T}\,)$ produces nonnegligible kinetic-energy dissipation with respect to all of the other schemes, which enforce the discretization of the pressure flux as $\overline{p}\,$. Fig. 3(c) shows the evolution of global internal energy, with an analogous production for the Gouasmi scheme due to the transfer of energy allowed by the pressure-work terms. The effect of the spatially integrated pressure work contribution is shown in Fig. 3(d), where the integrated time derivative of kinetic energy is shown as representative of the nonconservative pressure term in the discrete kinetic energy evolution equation, since the convective term is in locally conservative form. The nonzero average displayed by the Gouasmi scheme reflects its failure in keeping the global kinetic energy constant.
@@
An additional difference between the novel and previous formulations is seen in the fluctuations of the thermodynamic quantities of the density $\rho^{\prime}_{\mathrm{rms}}$ and temperature $T^{\prime}_{\mathrm{rms}}$, reported in Fig. 4, in which $\phi^{\prime}_{\mathrm{rms}}=\sqrt{\frac{1}{N_{p}}\sum_{i=1}^{N}(\phi_{i}-\phi_{\mu})^{2}}$ and $\phi_{\mu}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\phi_{i}$ is the average value of $\phi$ in the domain. In fact, as it has been already shown for the JP scheme ([12, 18]), it seems that energy fluxes whose pressure term is not built as in (2) (i.e. JP and Gouasmi formulations) produce fluctuations which do not reach a steady-state after the thermalization of the flow field, which is the expected result for homogeneous isotropic turbulence calculations [56].
4.3 Sod shock tube test
The EC schemes derived in the previous sections are constructed under the assumption of smooth solutions. However, compressible flows of practical interest typically involve discontinuities such as shocks and contact surfaces. In these regimes, purely non-dissipative discretizations are not appropriate: although they preserve a discrete entropy balance by construction, they do not enforce the entropy inequality required by the second law and generally produce spurious oscillations in under-resolved shocked regions.
A common strategy to address this issue is to employ a shock sensor and locally switch to a more dissipative discretization, such as WENO. This approach is followed, for instance, in [55], and entropy-stable (ES) variants of WENO have also been developed [57]. In the present work, we adopt an alternative and widely used strategy: rather than switching schemes, we augment the EC flux with a suitable dissipation term so as to obtain an ES formulation.
To recover physically admissible solutions, the semi-discrete scheme must satisfy a discrete entropy inequality instead of an exact entropy conservation property. This can be achieved by adding a consistent dissipative contribution to the EC numerical flux. Here, we consider a Rusanov (local Lax–Friedrichs, LLF) regularization, leading to the ES flux
$\mathcal{F}^{ES}(\mathbf{u}_{i},\mathbf{u}_{i+1})=\mathcal{F}^{EC}(\mathbf{u}_{i},\mathbf{u}_{i+1})-\frac{1}{2}\,\lambda_{\max}\left(\mathbf{u}_{i+1}-\mathbf{u}_{i}\right),$
where $\mathcal{F}^{EC}$ denotes the baseline entropy-conservative flux and $\lambda_{\max}$ is the maximum local signal speed,
$\lambda_{\max}=\max\left(|u_{i}|+c_{i},\;|u_{i+1}|+c_{i+1}\right),$
with $c$ the thermodynamic speed of sound of the TP gas model. This dissipation yields an ES scheme in the sense that the resulting formulation satisfies the correct entropy inequality at the semi-discrete level (see, e.g., [24]).
Figure 5: One-dimensional Sod shock tube problem. Results obtained with the EC-TP scheme augmented with a LLF dissipation term for different numbers of spatial discretization points $N_{x}$.
Although the LLF term is robust and straightforward to implement, it may introduce excessive diffusion, in particular across contact discontinuities. To mitigate this effect, we modulate the dissipative term with a simple pressure-jump sensor [31],
$\Xi=\sqrt{\frac{|p_{i+1}-p_{i}|}{p_{i+1}+p_{i}}},$
which scales the LLF contribution. The coefficient $\Xi$ is close to unity in the presence of strong pressure gradients (e.g. shocks) and small in smooth regions, thereby reducing unnecessary dissipation away from discontinuities. The resulting flux can be written as
$\mathcal{F}^{ES}=\mathcal{F}^{EC}-\frac{1}{2}\,\Xi\,\lambda_{\max}\left(\mathbf{u}_{i+1}-\mathbf{u}_{i}\right).$
To assess the performance of this ES extension in the presence of shocks, we consider the classical Sod shock tube problem. The one-dimensional computational domain $\Omega=[-1,1]$ is divided into two constant states by a diaphragm located at its midpoint, $x=0$. The initial left and right states are
$\begin{dcases}(\rho,u,p)=(1,0,1)\quad&\mathrm{for}\;\,x<0,\\ (\rho,u,p)=(0.125,0,0.1)\quad&\mathrm{for}\;\,x\geq 0.\end{dcases}$
At $t=0$ the diaphragm is removed, generating a right-moving shock, a left-moving rarefaction wave, and a contact discontinuity. The simulation is carried out up to $t/t_{\mathrm{ref}}=0.2$, where $t_{\mathrm{ref}}=L_{x}/c_{\mathrm{max}}$, with $c_{\mathrm{max}}=\max_{\Omega}(\sqrt{\gamma(T)RT})$ and $L_{x}=1$. CFL is equal to $0.1$, and it is based on the reference velocity $c_{\mathrm{max}}$.
Fig. 5 shows results for different numbers of discretization points $N_{x}=\{376,751,1501\}$ contained in the interval [1/2,1/2][-1/2,1/2]. Regarding spatial discretization, simulations have been carried out with the EC-TP formulation (Eqs. (2), (6) and (10)) for the smooth regions in its second-order accuracy version. The proposed strategy furnished excellent solutions which converge with increasing grid points. No visible oscillations are observed and the physical discontinuities are clearly highlighted.
@@ 5 Conclusions
Starting from the EC condition valid for a generic real gas, we demonstrated how the specific structure of the EoS for perfect gases can be exploited to eliminate singularities in the formulation of the numerical flux, thereby simplifying the resulting expressions. The only remaining singularity arises from the presence of logarithmic means, which can be addressed using locally the algorithm by Ismail and Roe [23] or by constructing a hierarchy of asymptotically entropy-conserving schemes thorough a Taylor series expansion, as in the case of calorically perfect gases. Importantly, this result is independent of the specific thermally perfect model used to relate internal energy and absolute temperature, supporting its use in a wide range of applications. As illustrative examples, we have considered both a polynomial-based model for the isochoric specific heat and the Rigid-Rotor Harmonic-Oscillator model, each leading to a different variant of the EC scheme.
@@
The results presented in this paper open several avenues for future development. For instance, further extensions could focus on deriving entropy-stable discretizations for the viscous terms as well, guaranteeing the correct entropy monotonicity, thus obtaining fully entropy-consistent spatial discretizations for arbitrary equations of state. Moreover, while the current formulation is restricted to nonreacting gas mixtures, the generality of the approach suggests potential extensions to reacting flows, including chemically reactive mixtures and multitemperature models.
@@ Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. We thank Dr. Ayaboe K. Edoh and Marco Artiano for reading an early draft of the manuscript and providing thoughtful feedback.